OpenStudy (anonymous):

solving systems of equations by elimination 7x+2y=24 8x+2y=30 explain if possible

OpenStudy (anonymous):

Just multiply the first equation by -1 and add it to the second equation... if it asks you to solve by elimination, it means you need to get rid of a variable in one of the equations...

OpenStudy (anonymous):

so your first equations becomes -7x-2y=-24 then add that too the second 8x+2y=30 ... that will then give you 1x = 6 then just substitute back in to either equation to get y

OpenStudy (anonymous):

8(6) + 2y = 30 2y = 30 - 48 2y = -18 y = -9

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

easy as pie. Use elimination method. 7x + 2y = 24 (1) 8x + 2y = 30 (2) Eliminate (1) and (2) 8x + 2y = 30 - 7x + 2y = 24 ------------------ x = 6 Substitute x = 6 into either (1) or (2) 8(6) + 2y = 30 2y = 30 - 48 y = -18/2 y = -9

OpenStudy (anonymous):

you can also check our answers. But with regards to the guy above me, the answer is right.

OpenStudy (hba):

@zonazoo You have provided the substitution method. @OrthodoxMan Try to interact with the user rather than providing solutions. The easy method to check your answer is to put the x and y values in the original question and check if L.H.S=R.H.S

OpenStudy (anonymous):

I get what you mean but some users learn by interpreting the solution rather than following a lesson.

OpenStudy (anonymous):

thank yall (:

OpenStudy (anonymous):

orthodox man helped alot

OpenStudy (hba):

@OrthodoxMan As a rule of thumb, we would prefer the askers to participate in the solution process instead of having the answerers doing all the work. thats not really our concern; we are aimed at being a study site, not a free answering service. We would rather promote learning the material as opposed to just handing over answers. This social contract is aimed at curbing people from using the site to cheat with ...

OpenStudy (anonymous):

and look who gets the best response @hba hahaha

OpenStudy (hba):

@zonazoo You provided the wrong method. The asker wanted to do it with Elimination method.

OpenStudy (anonymous):

@OrthodoxMan more than likely when you just type out the solution like that then they just go into a copy and paste format where they are not really thinking about the way it was solved. Also if you are staying with them for the entire question then they should be able to review the conversation just as they are able to review when you post just the solution

OpenStudy (anonymous):

@hba and what method are you suggesting I did... b/c what I did was no different than the other.

OpenStudy (anonymous):

@zonazoo you did elimination which is the same as addition but starts with the subtracting of one equation from the other instead, so you would not say to multiply by -1

OpenStudy (anonymous):

You're absolutely right, but this user really doesn't seem like she is doing a test at the moment. More like homework. However, I will keep it in mind for future questions.

OpenStudy (anonymous):

Thanks for the warning -.-

OpenStudy (anonymous):

lol @hba : this is elimination. Substitution is the other method and that deals with solving for one variable and putting it into the other equation. Medal for @zonazoo

OpenStudy (hba):

@Outkast3r09 I remember the first method as Substituion and the second as elimination.4 years is a long time dude :)

OpenStudy (anonymous):

@Outkast3r09 there are three methods for solving a system of equations 1. Substitution 2. Addition 3. Elimination I explained the difference between addition and elimination above

OpenStudy (hba):

I just checked both the methods and they are same :)

OpenStudy (anonymous):

oh i can name many other ways to solve this but the most used are substitution and elimination. In differential equations you solve these using matrices.

OpenStudy (anonymous):

I know my school system made sre we knew the difference between the two although it is an arbitrary difference

OpenStudy (anonymous):

I meant at this level if you want to start talk about RREF and graphing and so on we will be here all day listing all the ways to solve this question

OpenStudy (hba):

@OrthodoxMan Thanks a lot for understanding.

OpenStudy (anonymous):

I was talking about using a determinant but RREf works also

OpenStudy (anonymous):

addition and elimination are basically the same thing... really the only other basic way to solve these is to graph and find the intersection... other than that you would either just eliminate/add/subtract (whatever you want to call it) or substitute.

OpenStudy (anonymous):

and i think i meant linear algebra also hahaha.. differential could find all possibilities lol

OpenStudy (anonymous):

and the matrix needs to be a square matrix to take a determinant

OpenStudy (anonymous):

Now in regards with my warning, there was no need to give me one. You were already consulting me about this matter that there was no need to take it to issuing me a warning, Warnings are only for obvious mis-behaviours. I have committed none of that..

OpenStudy (anonymous):

Again I did admit that it always seemed like there wasn't really a difference between the two but I remember them calling them different names based on if we were subtracting or adding the two equations

OpenStudy (anonymous):

but either way... i think we have all just taken this topic far beyond what it needs to... im sure we all understand what is said.

OpenStudy (anonymous):

while yall are arguing haha who wants to answer the nest one the same way -4+9y=9 x-3y=-6

OpenStudy (hba):

@fkey3191 No one is going to answer your question,we can help you though.

OpenStudy (anonymous):

@fkey3191 I am guessing that is suppose to be -4x + 9y = 9 x - 3y = -6 otherwise that is way to simple

OpenStudy (anonymous):

not for me

OpenStudy (anonymous):

Okay then do they say it has to be elimination?

OpenStudy (anonymous):

Well with respect to my warning, refer to the method I used to solve your previous one and apply your knowledge on solving systems of equations from lessons. Before i get banned, goodluck.

OpenStudy (hba):

@OrthodoxMan If you have been issued a warning and you feel like you should not be warned,go talk to a mod ( the one in purple ) .

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

because one equation is already down to one variable which means you could just solve for y and sub in the answer for y into the second to solve for x . . . I do not see how this requires elimination at all.

OpenStudy (anonymous):

BTW I meant way to simple to use elimination that may not have been clear originally

OpenStudy (anonymous):

okay (:'

OpenStudy (anonymous):

Well then my last thing to add before leaving is (╯°□°)╯︵ ┻━┻ Why do they want to over complicate things?

OpenStudy (anonymous):

i know he was doing what i needed and i learn by getting the answer then solving up to that soo that guy was wrong.

OpenStudy (hba):

@fkey3191 If you learnt from the first question,why don't you imply it in the second question ?

OpenStudy (anonymous):

because i have 8 more pages and im not a smarty pants like you...and im not trying to all of them

OpenStudy (hba):

@fkey3191 I am sorry no one is going to do your homework here.

OpenStudy (anonymous):

take your brain and shove it up your retrice

OpenStudy (anonymous):

butt

OpenStudy (hba):

@Callisto Please look into the matter.

OpenStudy (hba):

@fkey3191 Using inappropriate speech is not allowed here,you may be banned to do so. And as i mentioned earlier We are here to help you not provide answers

OpenStudy (anonymous):

i said butt wow

OpenStudy (hba):

Please read the CoC- http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

i think you can read it for me

OpenStudy (callisto):

For asker, as always, this is a site for learning, not for answer-seeking. So, please do ask the others to do all the work for you (or give you the answer). For the ones who answer the question, we don't encourage you to give answer without the involving the asker in solving the question, since the asker may not learn and just copy your work. That's not learning. Regarding the warning, just take it as an reminder. If you don't make the same mistake again, that'd be fine. Last but the least, please mind the language used here.

OpenStudy (callisto):

Oh, I miss one thing there. Please start a new post for a new question.

OpenStudy (anonymous):

thats how i learn but okay.

OpenStudy (anonymous):

As i said, not everyone learns the same way. You guys impose quite a strict condition on people who cling to different methods of learning.

OpenStudy (callisto):

Hmm.. Perhaps at least the asker can show that he/she has done something but got stuck in somewhere?

OpenStudy (anonymous):

Besides, I'm giving a full solution and not just the answer.

OpenStudy (hba):

There should be some interaction @OrthodoxMan .

OpenStudy (anonymous):

As you saw in this example, it clearly did not work. http://openstudy.com/study#/updates/50b1bc5de4b0e906b4a63903

OpenStudy (callisto):

The keywords here are ''without the involving the asker in solving the question''

OpenStudy (callisto):

In some cases, it might not work, but at least for solving equations, you can let the asker participating in solving the question.

OpenStudy (anonymous):

Reword your quote here, it's grammatically incorrect.

OpenStudy (callisto):

Sorry, English is not my first language. Excuse me for my poor grammar/English. But do you get the meaning?

OpenStudy (anonymous):

Not really, I value the structure of sentences as it allows me to comprehend what you're saying. I would understand your "meaning" through verbal communication. Seems like that won't happen.

OpenStudy (hba):

'' without involving the asker in solving the question''*

OpenStudy (hba):

Now it sounds grammatically correct and i hope you understand ?

OpenStudy (hba):

If you are still having problems , Feel free to post a Question in the Open study feedback group- http://openstudy.com/study#/groups/OpenStudy%20Feedback.

OpenStudy (anonymous):

ahh that has been defined countless times in this thread. Now considering the fact I acknowledge who is asking the question and their purpose (say if they are taking a test or exam at the moment), you cannot issue a warning until it has been justified. If my claims are not justified, your goal to turn this site into a study group is not realistic. Why don't you check other threads for a 7 day time span and tell me whether people are giving answers or giving a path of way to the answer.

OpenStudy (hba):

@OrthodoxMan As i said if you have any further queries please post your questions in the Open study feedback group.